### v1

@@ Abstract
Accurate molecular property prediction is essential in drug discovery and related fields. However, existing graph neural networks (GNNs) often struggle to simultaneously capture both local and global molecular structures. In this work, we propose a Multi-Level Fusion Graph Neural Network (MLFGNN) that integrates Graph Attention Networks and a novel Graph Transformer to jointly model local and global dependencies. In addition, we incorporate molecular fingerprints as a complementary modality and introduce a mechanism of interaction between attention to adaptively fuse information across representations. Extensive experiments on multiple benchmark datasets demonstrate that MLFGNN consistently outperforms state-of-the-art methods in both classification and regression tasks. Interpretability analysis further reveals that the model effectively captures task-relevant chemical patterns, supporting the usefulness of multi-level and multi-modal fusion in molecular representation learning.
@@ 1 Introduction
In recent years, deep learning (DL) has achieved remarkable success in a variety of domains, such as computer vision 7 and natural language processing 8, largely driven by improvements in model architectures and the availability of large-scale data. These advances have inspired a wave of applications in scientific and biomedical fields, including atomic charge prediction 9, molecular property modeling 10, chemical reaction forecasting 11, and molecular generation 12. In particular, deep learning has shown great promise in molecular property prediction by learning generalizable representations directly from raw molecular data, thereby overcoming the limitations of handcrafted descriptors.
Among deep learning methods, graph neural networks (GNNs) have become the dominant paradigm for molecular modeling due to their ability to operate directly on molecular graph structures, where atoms are represented as nodes and bonds as edges. A variety of GNN architectures have been proposed for this task 13, 14, 15, 16. Among them, Graph Attention Networks (GATs) leverage attention mechanisms to adaptively weigh neighboring node features, making them effective for capturing local structural information 17, 18. In contrast, Graph Transformers incorporate global self-attention, enabling each node to interact with all others regardless of topological distance, which facilitates modeling global dependencies 19, 20, 21. Despite recent progress, existing models often fall short in addressing key challenges in molecular representation learning. For example, the well-known activity cliff phenomenon—where minor structural changes, such as the replacement of a single atom or functional group, can lead to dramatic shifts in biological activity—requires models that can simultaneously capture fine-grained local variations and global structural dependencies. Local-only models may overlook long-range effects, while global-only models risk missing chemically subtle but critical substructures. Moreover, feature representations such as molecular fingerprints, while chemically informative, may contain irrelevant or redundant elements that reduce prediction robustness. Compounding these challenges, different downstream tasks (e.g., solubility estimation vs. toxicity classification) often require attention to different aspects of molecular structure. These issues underscore the need for a unified framework capable of flexibly integrating diverse structural cues and adapting its focus based on task-specific demands. To address this issue, we design a multi-level fusion strategy that integrates both local and global information within the graph representation. Rather than statically combining the outputs of different modules 22, we introduce a learnable mechanism that adaptively balances their contributions based on task-specific needs.
In addition to structural information from molecular graphs, incorporating complementary modalities has shown promise for further enhancing predictive performance. Some approaches integrate molecular fingerprints 23, 24, motif-level features 25, 26, sequence representations 27, 3D conformational data 28, or even multiple sources simultaneously 29. Among these, molecular fingerprints are particularly valuable because they encode chemically meaningful substructures—such as functional groups, ring systems, or pharmacophores that are often directly linked to molecular activity. In comparison, other features may lack interpretability, contain redundant information, or introduce noise that is irrelevant to downstream tasks. Therefore, we choose molecular fingerprints as the complementary modality to be fused with graph-based representations. However, fingerprint vectors can also include irrelevant or noisy features that may harm predictive performance. Prior studies have largely overlooked this issue 23, 24. To overcome it, we introduce a cross-attention mechanism that enables the model to adaptively aggregate and filter features from both modalities, ensuring that only task-relevant information contributes to the final molecular representation.
@@
In this study, we propose a Multi-Level Fusion Neural Network for molecular property prediction, which integrates both intra-graph (local–global) and inter-modal (graph–fingerprint) fusion. The graph branch combines local and global structural cues through the integration of GATs and a newly designed Graph Transformer. The fingerprint branch introduces domain-informed substructure features. A final cross-attention layer performs adaptive aggregation and feature selection across modalities. Extensive experiments on multiple benchmark datasets demonstrate the effectiveness, interpretability, and generalizability of our approach.
@@ 2.2 The architecture of MLFGNN
3:$M_{i}=Morgan(S),P_{i}=PubChem(S),E_{i}=PharmaERG(S)$
4:$U_{i}=[M_{i}\text{ }||\text{ }P_{i}\text{ }||\text{ }E_{i}]$
5:$FP_{i}=MLP(U_{i})$
@@
Figure 1: The Multi-Level Fusion Graph Neural Network Architecture (A) The overview of model. (B) The architecture Mixed-information Layer.
@@ 2.3 Molecular fingerprints network
To capture comprehensive molecular-level information, we constructed a composite fingerprint representation by concatenating three widely used molecular descriptors: Morgan fingerprint 30 $M_{i}$, PubChem Fingerprint 31 $P_{i}$, and Pharmacophore ErG Fingerprint 32 $E_{i}$. Each fingerprint captures different aspects of the molecular structure and properties. Through the smiles $x_{i}$ of the molecules, we can get the corresponding molecular fingerprints. The Morgan Fingerprint encodes circular substructures, PubChem Fingerprint incorporates a variety of predefined structural keys, and ErG Fingerprint reflects pharmacophore features relevant for ligand-receptor interactions. By combining these fingerprints, we form a unified vector representation $U_{i}$ that encompasses a broad spectrum of molecular characteristics,
$U_{i}=[M_{i}\text{ }||\text{ }P_{i}\text{ }||\text{ }E_{i}].$ (1)
@@
This embedding is subsequently used in down stream prediction tasks within out neural network framework so that it can get global information and can choose some information that is useful for predicting the molecular property.
@@ 2.4 Molecular featurization
To featurize the nodes and edges in the molecular graph, we encode them into vectors derived from a set of well-established atomic and bond characteristics. In this work, we employ 16 types of atomic features and 4 types of bond features to effectively capture the chemical identity of atoms and their local chemical environment (see Table 1). All features are encoded using one-hot vectors, except for atomic mass and formal charge, which are preserved as continuous numerical values. For atom features, since hydrogen atoms are not explicitly represented in SMILES strings, the implicit valence of atoms is used as a proxy to infer the presence of bonded hydrogens thereby preserving chemical completeness. Additionally, we incorporate whether an atom functions as a hydrogen bond donor or acceptor, and whether it is part of an acidic or basic function group, as these properties are critical for molecular interactions and biological activity. For bond features, stereo-chemical information is also included, as it reflects the spatial arrangement of atoms, which can significantly influence molecular properties and interactions. As a result, for each molecular graph $G_{i}=\{\mathcal{V}_{i},\mathcal{E}_{i}\}$, each node $v\in\mathcal{V}_{i}$ is represented by a 57-dimensional feature vector $A_{v}$, and each edge $e_{vu}\in\mathcal{E}_{i}$ is encoded with a 13-dimensional feature vector $B_{vu}$.
@@ 2.5 Molecular graph Representation
Building upon the Attentive FP framework 33, which employs graph attention networks (GAT17), our approach utilizes a similar graph attention mechanism to encode molecular graphs effectively. However, unlike Attentive FP, which leverages deep multi-layer attention to capture global dependencies across the molecule, our design intentionally employs a shallower attention stack. This enables the model to focus on local substructure patterns, which are often more informative for certain molecular properties(e.g., functional groups or local motifs). In particular, we apply only a few attention layers to preserve locality and reduce over-smoothing, allowing the model to capture neighborhood-level information more effectively. To formally describe our Graph Attention Layer, consider a molecular graph $G_{i}=\{\mathcal{V}_{i},\mathcal{E}_{i}\}$, where $\mathcal{V}_{i}$ and $\mathcal{E}_{i}$ denote the sets of atoms (nodes) and bonds(edges), respectively. The initial feature representations for atoms $A_{v}$ and bonds $B_{vu}$ are constructed using a set of domain-specific chemical descriptors, capturing both intrinsic and relational chemical properties.
@@
This neighbor representation is further processed through a fully connected layer with ReLU activation function
@@
To enhance the expressive capacity of the model, we update the node embedding using a GRU34, which integrates the aggregated neighborhood information with the previous hidden state:
@@
While GAT captures local motifs, long-range dependencies require a global mechanism. We adopt a Transformer-based architecture enhanced with structural priors and the detailed structure of the new Transformer architecture is illustrated in Figure 2. Specifically, we integrate the graph’s normalized adjacency matrix $A$ into the self-attention weights:
@@
where $Q_{i},K_{i},V_{i}$ are query, key, and value matrices; $\lambda_{a},\lambda_{b}$ are learnable scalars controlling the balance between learned attention and explicit graph structure.
Additionally, we replace the conventional Layer Normalization used in Transformer blocks with a novel Dynamic Tanh(DyT) 35 layer. DyT is a recently proposed alternative to traditional normalization layers, introduced as a lightweight and effective element-wise nonlinearity. Unlike Layer Normalization, which requires computing per-token statistics to rescale activations, DyT operates via a simple learnable squashing function defined as:
@@
where $\alpha$ is a learnable scalar that dynamically scales the input, and $\gamma,\beta$ are learnable affine parameters similar to those in normalization layers. This formulation preserves the S-shaped activation squashing effect observed in LayerNorm outputs, particularly for extreme activation values, while significantly reducing computational overhead.
@@
To integrate these two perspectives, we introduce a Mixed Information Layer that adaptively fuses the local and global features. First, we perform inter-layer aggregation over the outputs of all GAT layers. This is motivated by the observation that different GAT layers capture different neighborhood sizes (receptive fields), and aggregating them provides a more expressive local representation. The average over all GAT layers is computed as:
@@
This equation ensures an adaptive balance between local substructure sensitivity ($F_{A}$) and global topological awareness ($F_{T}$), allowing the model to modulate emphasis depending on the molecular property being predicted. For instance, properties like reactivity may benefit more from local patterns, whereas solubility may require more global context. To convert node-level features into a single molecular representation, we introduce a Virtual Super Node $H_{v}$, which serves as a global summarization anchor. It is initialized as the sum of all atomic embeddings from the fused representation:
@@
where $[H_{v}||H_{i}]$ denotes concatenation of the global super node and atomic embedding, $W_{1}\in\mathbb{R}^{d\times 2d}$ is a learnable weight matrix, and $e_{vi}\in\mathbb{R}$ is an unnormalized scalar attention score reflecting the interaction between the $i$-th atom and the global context. These scores are normalized via softmax to obtain attention weights:
@@
where $W_{2}\in\mathbb{R}^{d\times d}$ is a learnable transformation matrix, and $elu$ provides smooth activation to stabilize training and enhance expressiveness. $C_{v}\in\mathbb{R}^{d}$ captures a weighted summary of atomic features, focused on the most informative atoms. Finally, we update the virtual node $H_{v}$ using a GRU to integrate this attention-refined context:
@@
The Virtual Super Node $F_{v}$ serves as a global, task-adaptive representation of the entire molecule. It is subsequently passed into a MLP for downstream prediction tasks. Importantly, the introduction of this virtual node is not merely for summarization, but plays a crucial role in further refining the fused local-global information obtained from Adaptive Mixture Layer. By attending to most relevant atomic embeddings via an attention mechanism and incorporating this context through a GRU-based update, the virtual node effectively filters and distills the most predictive aspects of the mixed features. This process enhances the model’s capacity to generate expressive and discriminative molecular representations, ultimately improving its performance on various prediction tasks
Figure 2: The details of Graph Transformer Layer.
@@ 3.1 Benchmark datasets
In this study, we evaluated the performance of our model on 10 publicly available data sets from Wu et al36. These include five classification datasets: BACE, BBBP, Tox21, SIDER, and ClinTox, as well as five regression datasets including ESOL, FreeSolv, PDBbind-C, PDBbind-R and Lipophilicity. The classification tasks primarily focus on the prediction of physiology or biophysics, with the evaluation based on the ROC-AUC metric. The regression tasks target physicochemical properties and are evaluated using RMSE. These datasets span diverse molecular properties and cover a broad spectrum of molecular structures, providing a comprehensive testbed for assessing the generalization capability of our model. More information about these datasets can be found in Table 2.

### v2

@@ Abstract
Accurate prediction of molecular properties is essential in drug discovery and related fields. However, existing graph neural networks (GNNs) often struggle to simultaneously capture both local and global molecular structures. In this work, we propose a Multi-Level Fusion Graph Neural Network (MLFGNN) that integrates Graph Attention Networks and a novel Graph Transformer to jointly model local and global dependencies. In addition, we incorporate molecular fingerprints as a complementary modality and introduce a mechanism of interaction between attention to adaptively fuse information across representations. Extensive experiments on multiple benchmark datasets demonstrate that MLFGNN consistently outperforms state-of-the-art methods in both classification and regression tasks. Interpretability analysis further reveals that the model effectively captures task-relevant chemical patterns, supporting the usefulness of multi-level and multi-modal fusion in molecular representation learning.
@@ 1 Introduction
In recent years, deep learning (DL) has achieved remarkable success in a variety of domains, such as computer vision 7 and natural language processing 8, largely driven by improvements in model architectures and the availability of large-scale data. These advances have inspired a wave of applications in scientific and biomedical fields, including atomic charge prediction 9, molecular property modeling 10, chemical reaction forecasting 11, molecular generation 12 and drug property prediction13, 14, 15, 16. In particular, deep learning has shown great promise in molecular property prediction by learning generalizable representations directly from raw molecular data, thereby overcoming the limitations of handcrafted descriptors.
Among deep learning methods, graph neural networks (GNNs) have become the dominant paradigm for molecular modeling due to their ability to operate directly on molecular graph structures, where atoms are represented as nodes and bonds as edges. A variety of GNN architectures have been proposed for this task 17, 18, 19, 20. However, many of these architectures struggle to effectively integrate both local chemical environments and global molecular context, limiting their capacity to capture complex structure-property relationships. One prominent example of such complexity is the activity cliff phenomenon, in which small structural modifications, such as the addition, removal, or substitution of a single atom or function group, can result in disproportionately large changes in biological activity. Accurately modeling activity cliffs requires molecular representations that are both locally sensitive and globally aware, which many existing GNN-based models fail to achieve. To address this limitation, many researches have explored pretraining strategies to enhance the ability of GNNs to capture both local and global molecular information21, 22, 23, 24. However, pretraining approaches may suffer from negative transfer, where the knowledge learned during pretraining adversely affects performance on downstream tasks, thereby limiting their effectiveness. In light of these limitations, an alternative line of research focuses on architectural innovations that can inherently balance local and global molecular information without relying heavily on pretraining.
In response to these challenges, we design a unified framework that simultaneously address two key aspects of molecular representation: the fusion of local and global structural information, and the incorporation of chemically meaningful features from complementary modalities. While GATs25, 26 and Graph Transformers27, 28, 29 have shown promise individually, their combination offers a more comprehensive view of molecular graphs by capturing both neighborhood-level patterns and long-range dependencies.
To further enhance molecular representation, we incorporate molecular fingerprints as an additional modality, motivated by their ability to encode domain knowledge in the form of predefined substructures, such as functional groups, ring systems, and pharmacophores30, 31. These features offer strong chemical priors that are particularly useful for capturing subtle structural motifs, especially in scenarios like activity cliffs where small local changes can lead to significant functional differences. Moreover, integrating fingerprints with graph-based representations can compensate for each other’s limitations: while molecular graphs provide flexible and learnable topology-aware embeddings, fingerprints contribute fixed, interpretable descriptors that may be more robust to noise in graph structure or representation sparsity. Previous studies have demonstrated the benefits of such multi-modal integration32, 33, but have largely overlooked the importance of selectively filtering task-irrelevant or redundant fingerprint features, which may degrade performance if naively combined30, 31 . To address this, we introduce a cross-attention mechanism that adaptively aggregates and filters information from both molecular graphs and fingerprints, ensuring that only task-relevant signals contribute to the final representation.
@@
Building upon these insights, we propose the Multi-Level Fusion Graph Neural Network(MLFGNN) for molecular property prediction, which performs both intra-graph(local-global) and inter-modal(graph-fingerprint) fusion. The graph branch integrates GATs and a newly designed graph Transformer to effectively capture hierarchical structural information within molecular graphs. In parallel, the fingerprint branch incorporates domain-informed substructure features that provide complementary chemical priors. To harmonize these two modalities, a final cross-attention layer adaptive aggregates and filters their representations, allowing the model to focus on task-relevant features. Extensive experiments on multiple benchmark datasets demonstrate that MLFGNN achieves strong predictive performance while offering improved interpretability and generalization across diverse molecular tasks.
Our contributions can be summarized as follows:
  • Introduction of a novel Graph Transformer architecture integrated with Graph Attention Networks(GATs), enabling more effective fusion of local and global molecular representations.
  • A cross-modal fusion strategy combining graph-based and fingerprint-based features using adaptive attention.
  • Empirical validation through ablation and visualization studies demonstrates the model’s generalizability and interpretability, highlighting its focus on chemically relevant substructures.
@@ 2.2 The architecture of MLFGNN
[…]
@@
Figure 1: The Multi-Level Fusion Graph Neural Network Architecture (A) The overview of model. (B) The architecture Graph Tansformer Layer. (C)The details of Molecule Attention Layer (D)The Message Passing process of Graph Attention Layer.
@@ 2.3 Molecular fingerprints network
To capture comprehensive molecular-level information, we constructed a composite fingerprint representation by concatenating three widely used molecular descriptors: Morgan fingerprint 34 $M_{i}$, PubChem Fingerprint 35 $P_{i}$, and Pharmacophore ErG Fingerprint 36 $E_{i}$. The concatenation of Morgan, PubChem, and Pharmacophore ErG fingerprints provides a complementary representation of molecular structures by capturing distinct yet synergistic aspects of chemical information. Specifically, the Morgan fingerprint encodes circular substructures and topological environments, the PubChem fingerprint reflects the presence of predefined substructure keys and function groups, while the Pharmacophore ErG fingerprint emphasizes 3D pharmacophoric patterns and spatial relationships relevant to bioactivity, thereby collectively enhancing the molecular representation for predictive modeling. Through the smiles $x_{i}$ of the molecules, we can get the corresponding molecular fingerprints. The Morgan Fingerprint encodes circular substructures, PubChem Fingerprint incorporates a variety of predefined structural keys, and ErG Fingerprint reflects pharmacophore features relevant for ligand-receptor interactions. By combining these fingerprints, we form a unified vector representation $U_{i}$ that encompasses a broad spectrum of molecular characteristics,
[…]
@@
This embedding is subsequently used in downstream prediction tasks within our neural network framework so that it can get global information and choose some information that is useful for predicting the molecular property.
@@ 2.4 Molecular featurization
To featurize the nodes and edges in the molecular graph, we encode them into vectors derived from a set of well-established atomic and bond characteristics. In this work, we employ 16 types of atomic features and 4 types of bond features to effectively capture the chemical identity of atoms and their local chemical environment (see Table 1). All features are encoded using one-hot vectors, except for the atomic mass and formal charge, which are preserved as continuous numerical values. For atom features, since hydrogen atoms are not explicitly represented in SMILES strings, the implicit valence of atoms is used as a proxy to infer the presence of bonded hydrogens thereby preserving chemical completeness. Additionally, we incorporate whether an atom functions as a hydrogen bond donor or acceptor, and whether it is part of an acidic or basic function group, as these properties are critical for molecular interactions and biological activity. For bond features, stereo-chemical information is also included, as it reflects the spatial arrangement of atoms, which can significantly influence molecular properties and interactions. As a result, for each molecular graph $G_{i}=\{\mathcal{V}_{i},\mathcal{E}_{i}\}$, each node $v\in\mathcal{V}_{i}$ is represented by a 57-dimensional feature vector $A_{v}$, and each edge $e_{vu}\in\mathcal{E}_{i}$ is encoded with a 13-dimensional feature vector $B_{vu}$.
@@ 2.5 Molecular graph Representation
Building on the Attentive FP framework 37, which employs graph attention networks (GAT25), our approach utilizes a similar graph attention mechanism to effectively encode molecular graphs. However, unlike Attentive FP, which leverages deep multi-layer attention to capture global dependencies across the molecule, our design intentionally employs a shallower attention stack. This enables the model to focus on local substructure patterns, which are often more informative for certain molecular properties(e.g., functional groups or local motifs). In particular, we apply only a few attention layers to preserve locality and reduce over-smoothing, allowing the model to capture neighborhood-level information more effectively. To formally describe our Graph Attention Layer, consider a molecular graph $G_{i}=\{\mathcal{V}_{i},\mathcal{E}_{i}\}$, where $\mathcal{V}_{i}$ and $\mathcal{E}_{i}$ denote the sets of atoms (nodes) and bonds(edges), respectively. The initial feature representations for atoms $A_{v}$ and bonds $B_{vu}$ are constructed using a set of domain-specific chemical descriptors, capturing both intrinsic and relational chemical properties.
@@
This neighbor representation is further processed through a fully connected layer with ReLU activation function:
@@
To enhance the expressive capacity of the model, we update the node embedding using a GRU38, which integrates the aggregated neighborhood information with the previous hidden state:
@@
While GAT effectively captures local structural motifs through neighborhood attention, modeling long-range dependencies requires a more global mechanism. To address this, we adopt a Transformer-based architecture augmented with structural priors derived from the graph topology. The detailed structure of our modified Transformer is illustrated in Figure 1(B).Specifically, we incorporate the graph’s normalized adjacency matrix $A$ into the self-attention computation to inject explicit structural information. The attention operation is modified as follows:
@@
where $Q_{i},K_{i}$ and $V_{i}$ are query, key, and value matrices, respectively, and $\lambda_{a},\lambda_{b}$ are learnable scalars balancing the learned attention and the structural prior.
In this formulation, the softmax term provides global context by enabling interactions between all node pairs, while the adjacency matrix $A$ reinforces the known connections from the molecular graph. This design ensures that the model captures both long-range dependencies and local structural fidelity. In particular, for molecules with sparse or branched topologies, the learned attention offers flexible contextualization, while $A$ maintains connectivity-based inductive bias. For densely connected molecules, $A$ further stabilizes attention by anchoring it to chemically meaningful neighborhoods.
Additionally, we replace the conventional Layer Normalization used in Transformer blocks with a novel Dynamic Tanh(DyT) 39 layer. DyT is a recently proposed alternative to traditional normalization layers, introduced as a lightweight and effective element-wise nonlinearity. Unlike Layer Normalization, which requires computing per-token statistics to rescale activations, DyT operates via a simple learnable squashing function defined as:
@@
where $\alpha$ is a learnable scalar that dynamically scales the input, and $\gamma,\beta$ are learnable affine parameters similar to those in normalization layers. This formulation preserves the S-shaped activation squashing effect observed in LayerNorm outputs, particularly for extreme activation values, while significantly reducing computational overhead. We have done relevant experiments and put them in the Supporting Information.
@@
To integrate these two perspectives, we introduce a Mixed Information Layer that adaptively fuses local and global features. First, we perform inter-layer aggregation over the outputs of all GAT layers. This is motivated by the observation that different GAT layers capture different neighborhood sizes (receptive fields), and aggregating them provides a more expressive local representation. The average over all GAT layers is computed as:
@@
This equation ensures an adaptive balance between local substructure sensitivity ($F_{A}$) and global topological awareness ($F_{T}$), allowing the model to modulate emphasis depending on the molecular property being predicted. For instance, properties like reactivity may benefit more from local patterns, whereas solubility may require more global contexts. To convert node-level features into a single molecular representation, we introduce a Virtual Super Node $H_{v}$, which serves as a global summarization anchor. It is initialized as the sum of all atomic embeddings from the fused representation:
@@
where $H_{v}||H_{i}]$ denotes concatenation of the global super node and atomic embedding, $W_{1}\in\mathbb{R}^{d\times 2d}$ is a learnable weight matrix, and $e_{vi}\in\mathbb{R}$ is an unnormalized scalar attention score reflecting the interaction between the $i$-th atom and the global context. These scores are normalized via softmax to obtain attention weights:
@@
where $W_{2}\in\mathbb{R}^{d\times d}$ is a learnable transformation matrix, and eluelu provides smooth activation to stabilize training and enhance expressiveness. $C_{v}\in\mathbb{R}^{d}$ captures a weighted summary of atomic features, focused on the most informative atoms. Finally, we update the virtual node $H_{v}$ using a GRU to integrate this attention-refined context:
@@
The Virtual Super Node $F_{v}$ serves as a global, task-adaptive representation of the entire molecule. It is subsequently passed into a MLP for downstream prediction tasks. Importantly, the introduction of this virtual node is not merely for summarization, but plays a crucial role in further refining the fused local-global information obtained from Adaptive Mixture Layer. By attending to most relevant atomic embeddings via an attention mechanism and incorporating this context through a GRU-based update, the virtual node effectively filters and distills the most predictive aspects of the mixed features. This process enhances the model’s capacity to generate expressive and discriminative molecular representations, ultimately improving its performance on various prediction tasks.
@@ 3.1 Benchmark datasets
In this study, we evaluated the performance of our model on 10 publicly available data sets from Wu et al40. These include five classification datasets: BACE, BBBP, Tox21, SIDER, and ClinTox, as well as five regression datasets including ESOL, FreeSolv, PDBbind-C, PDBbind-R and Lipophilicity. The classification tasks primarily focus on the prediction of physiology or biophysics, with the evaluation based on the ROC-AUC metric. The regression tasks target physicochemical properties and are evaluated using RMSE. These datasets span diverse molecular properties and cover a broad spectrum of molecular structures, providing a comprehensive testbed for assessing the generalization capability of our model. More information about these datasets can be found in Table 2.